OpenStudy (anonymous):

An airplane starts from rest and accelerates at 5.99 m/s2. What is its speed at the end of a 500-m runway?

OpenStudy (anonymous):

V(final)^2 = V(initial)^2 +2*a*x you can use this equation.

OpenStudy (anonymous):

^ vf^2=vi^2+2ad vf^2=2ad

OpenStudy (anonymous):

Supposing it travels in a straight line, one could use the kinematic equation for uniformly accelerated motion. \[FinalVelocity^{2} = InitialVelocity^{2} + 2 * Acceleration * Distance\] Which, provided that you start from rest, and travel 500m, becomes: \[FinalVelocity^{2} = 2 * 5.99 \frac{ m }{ s^{2} } * 500 m\] \[FinalVelocity^{2} = 5990 \frac{ m^{2} }{ s^{2} }\] \[FinalVelocity = \begin{matrix}+ \\ -\end{matrix} 77.40 \frac{ m }{ s }\] We assume that the plane does not fly backwards, so: \[FinalVelocity = + 77.40 \frac{ m }{ s }\]

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."